# thanks: Corresponding author

### v1

@@ III Results and Discussion
It is clear from Table 1 and Figure 2 (b) that the convergence of the lowest-energy ${}^{1}A_{g}(D_{2\text{h}})$ potentials resulting from the CIPSI-driven CC($P$;$Q$) calculations toward their CCSDT parent, including the challenging barrier region, with the CIPSI wave function termination parameter $N_{\text{det(in)}}$, with the number of determinants in the final Hamiltonian diagonalization space used to determine $|\Psi^{(\text{CIPSI})}\rangle$ [$N_{\text{det(out)}}$], and with the fraction of triply excited determinants in the $P$ space captured by CIPSI is very fast, but one cannot say the same about the uncorrected CC($P$) energies. As shown in Table 1 and Figure 2 (a), and in line with the formal analysis in Section II, the CC($P$) energies improve the CCSD results and converge toward CCSDT, but they do it at a much slower rate than their $\delta(\mbox{$P$;$Q$})$-corrected CC($P$;$Q$) counterparts. For instance, the CIPSI-driven CC($P$) computations for the lowest singlet state of cyclobutadiene using $N_{\text{det(in)}}=100,000$ reduce the 26.827, 27.964, 29.667, 32.473, 37.662, and 47.979 millihartree errors relative to CCSDT obtained at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 with CCSD and the associated NPE value of 21.152 millihartree to 22.183, 23.029, 23.947, 30.198, 21.842, 32.125, and 10.283 millihartree, respectively. This should be compared to the much smaller error and NPE values characterizing the corresponding CC($P$;$Q$) calculations, which are 0.431–3.539 and 3.108 millihartree, respectively. The analogous CC($P$) calculations using $N_{\text{det(in)}}=250,000$, where the errors characterizing the $\delta(\mbox{$P$;$Q$})$-corrected CC($P$;$Q$) energies in the entire $\lambda=0$–1 region and the overall NPE relative to CCSDT are already at the level of 0.2–0.5 millihartree, produce the 17.137–25.279 millihartree errors and the NPE of 8.142 millihartree. Even with the largest $N_{\text{det(in)}}$ value considered in this study, of $10,000,000$, the 7.252–10.238 millihartree differences between the CC($P$) and CCSDT energies of the lowest ${}^{1}A_{g}(D_{2\text{h}})$ state of cyclobutadiene in the $\lambda=0$–1 region and the NPE of 2.986 millihartree that characterizes the resulting CC($P$) potential relative to its CCSDT parent remain. All of this implies that while relaxing $T_{1}$ and $T_{2}$ amplitudes in the presence of the leading triples is important, correcting the CC($P$) energies for the remaining $T_{3}$ effects, which the CC($P$) computations using the $P$ spaces generated with the help of CIPSI do not describe, is critical to reach submillihartree accuracy levels relative to CCSDT with small fractions of triples in these spaces. We observed a similar behavior in the semi-stochastic, CIQMC- and CCMC-driven,Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022) and adaptiveGururangan and Piecuch (2023) CC($P$) and CC($P$;$Q$) calculations, although based on the numerical evidence that we have generated to date, the CIPSI-driven CC($P$;$Q$) methodology investigated in this study and its recently formulated adaptive analog seem to be more effective in converging the target CC (in most of our work to date, CCSDT) energetics than their semi-stochastic counterparts. This suggests that the sequences of Hamiltonian diagonalizations utilized in the CIPSI-driven CC($P$) and CC($P$;$Q$) computations and the moment expansions defining the $\delta(\mbox{$P$;$Q$})$ corrections that are used to construct excitation spaces in the adaptive CC($P$) and CC($P$;$Q$) runs are more efficient in identifying the leading higher–than–doubly excited determinants for inclusion of the underlying $P$ spaces than the CIQMC/CCMC wave function propagations, although this topic needs to be explored further and we will return to it in the future.
@@
As shown in Table 2 and Figure 2 (c) and (d), many of the above observations apply to the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations for the lowest-energy ${}^{3}B_{1g}(D_{\rm 2h})$ potential, but, given the fact that the CR-CC(2,3) approach is already very accurate in this case, producing errors relative to CCSDT that in absolute value do not exceed 60 microhartree, the CC($P$;$Q$) computations using $N_{\text{det(in)}}>1$ offer no obvious advantages over CR-CC(2,3). Nonetheless, it is reassuring that, in analogy to the $A_{g}(D_{\rm 2h})$-symmetric singlet ground state, the differences between the energies of the ${}^{3}B_{1g}(D_{\rm 2h})$ state obtained in the CC($P$) computations using CIPSI Hamiltonian diagonalizations to create lists of the leading triply excited determinants for inclusion in the underlying $P$ spaces and their CCSDT counterparts decrease as $N_{\text{det(in)}}$ increases, independent of $\lambda$. It is also encouraging that the $\delta(\mbox{$P$;$Q$})$ corrections are as effective in bringing the CC($P$) energies to a virtually perfect agreement with the parent CCSDT data as in the case of the CIPSI-driven CC($P$;$Q$) calculations for the lowest ${}^{1}A_{g}(D_{2\text{h}})$ potential using $N_{\text{det(in)}}\geq 250,000$. By inspecting the CC($P$;$Q$) column in Table 2, one may get the impression that the incorporation of the triply excited determinants identified by the CIPSI runs using increasingly large $N_{\text{det(in)}}$ values in the preceding CC($P$) steps worsens the CR-CC(2,3) results for the lowest ${}^{3}B_{1g}(D_{\rm 2h})$ state, which correspond to $N_{\text{det(in)}}=1$, but reading Table 2 in this way would be misleading. Indeed, in single-reference situations, such as that created by the lowest ${}^{3}B_{1g}(D_{\rm 2h})$ state of cyclobutadiene, where many-electron correlation effects are essentially only dynamical, the triples correction of CR-CC(2,3) [similarly to CCSD(T)] often overshoots the parent CCSDT energies (slightly). Once one starts adding triply excited determinants to the $P$ space, the CC($P$;$Q$) energies initially go up, becoming upper bounds to their CCSDT counterparts, but when the fraction of triples in the $P$ space is large enough, the differences between the CC($P$;$Q$) and CCSDT energies decrease, steadily approaching 0. We see some of this behavior in Table 2, but we have to keep in mind that the $P$ spaces used in our CIPSI-driven CC($P$;$Q$) calculations use small fractions of triples, so that the residual, $\sim$0.1 millihartree, errors relative to CCSDT remain. What is most important here is that the CC($P$;$Q$) computations for the lowest triplet potential of cyclobutadiene using $N_{\text{det(in)}}>1$ reported in Table 2 and Figure 2 (d) do not substantially alter the already excellent CR-CC(2,3) energetics. This allows us to conclude that the coupling of the lower-rank $T_{1}$ and $T_{2}$ clusters with their higher-rank $T_{3}$ counterpart is negligible in this case and the relaxation of the CCSD values of $T_{1}$ and $T_{2}$ amplitudes by including some triples in the iterative CC($P$) steps is not necessary for obtaining high-accuracy CC($P$;$Q$) results.
@@
Having demonstrated the excellent performance of the CIPSI-driven CC($P$;$Q$) approach in accurately approximating the lowest-energy ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ potentials of cyclobutadiene along its automerization coordinate obtained with CCSDT, in the final part of our discussion, we turn to the CC($P$) and CC($P$;$Q$) singlet–triplet gaps and their dependence on $\lambda$ and $N_{\text{det(in)}}$ examined in Table 3 and Figure 2 (e) [the uncorrected CC($P$) energetics] and (f) [the CC($P$;$Q$) results]. As pointed out above, to obtain accurate $\Delta E_{\text{S--T}}$ values for cyclobutadiene in the vicinity of the barrier region, one has to balance significant nondynamical correlations associated with the multiconfigurational singlet state, which manifest themselves in massive $T_{3}$ clusters that are strongly coupled to the one- and two-body components of $T$, with predominantly dynamical correlations characterizing the lowest triplet state that result in the generally smaller $T_{3}$ contributions having minimal effect on $T_{1}$ and $T_{2}$ amplitudes. The singlet–triplet gap values reported in Table 3 and Figure 2 (e) and (f) clearly show that neither the CCSD approach nor the CR-CC(2,3) triples correction to CCSD, which are equivalent to the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations using $N_{\text{det(in)}}=1$, can do this. Both of these methods struggle with achieving a balanced description of the ${}^{1}A_{g}(D_{\rm 2h})$ and ${}^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene as $\lambda$ approaches 1, producing errors relative to CCSDT that are as large as 5.261 and 2.282 kcal/mol, respectively, at $\lambda=0.6$, where the CCSDT value of $\Delta E_{\text{S--T}}$ is $-10.295$ kcal/mol, 8.589 and 4.434 kcal/mol, respectively, at $\lambda=0.8$, where $\Delta E_{\text{S--T}}$ obtained with CCSDT is $-5.804$ kcal/mol, and 15.120 and 9.222 kcal/mol at $\lambda=1$, where the CCSDT result for $\Delta E_{\text{S--T}}$ is $-4.783$ kcal/mol. These large error values in the singlet–triplet gaps resulting from the CCSD and CR-CC(2,3) computations in the barrier region are a consequence of the dramatic increase in the magnitude of $T_{3}$ effects characterizing the ${}^{1}A_{g}(D_{\rm 2h})$ state and a rapidly deteriorating description of this state by both CCSD and CR-CC(2,3) as $\lambda\rightarrow 1$, seen in Table 1 and Figure 2 (a) and (b), as opposed to the nearly constant $T_{3}$ contributions and a virtually perfect agreement between the CR-CC(2,3) and CCSDT ${}^{3}B_{1g}(D_{\rm 2h})$ potentials in the entire $\lambda=0$–1 region shown in Table 2 and Figure 2 (d) [while quantitatively inaccurate, the shape of the ${}^{3}B_{1g}(D_{\rm 2h})$ potential obtained with CCSD, shown in Figure 2 (c), is qualitatively correct too]. In analogy to the previously discussed calculations for the lowest-energy ${}^{1}A_{g}(D_{\rm 2h})$ potential, in order to bring the above errors down, the input parameter $N_{\text{det(in)}}$, which controls the CIPSI diagonalization sequences preceding the CC($P$) and CC($P$;$Q$) computations, must be increased, so that the $P$ spaces used in these computations are augmented with the leading triply excited determinants, the CCSD values of $T_{1}$ and $T_{2}$ amplitudes, used in CR-CC(2,3), are properly relaxed, and the quality of the $\delta(\mbox{$P$;$Q$})$ corrections that capture the remaining $T_{3}$ effects improves.

### v2

@@ III Results and Discussion
It is clear from Table 1 and Figures 2 and 5 (b) that the convergence of the lowest-energy $^{1}A_{g}(D_{2\text{h}})$ potentials resulting from the CIPSI-driven CC($P$;$Q$) calculations toward their CCSDT parent, including the challenging barrier region, with the CIPSI wave function termination parameter $N_{\text{det(in)}}$, with the number of determinants in the final Hamiltonian diagonalization space used to determine $|\Psi^{(\text{CIPSI})}\rangle$ $N_{\text{det(out)}}$], and with the fraction of triply excited determinants in the $P$ space captured by CIPSI is very fast, but one cannot say the same about the uncorrected CC($P$) energies. As shown in Table 1 and Figures 2 and 5 (a), and in line with the formal analysis in Section II, the CC($P$) energies improve the CCSD results and converge toward CCSDT, but they do it at a much slower rate than their $\delta(\mbox{$P$;$Q$})$-corrected CC($P$;$Q$) counterparts. For instance, the CIPSI-driven CC(PP)/$c$-pVDZ computations for the lowest singlet state of cyclobutadiene using $N_{\text{det(in)}}=100,000$ reduce the 26.827, 27.964, 29.667, 32.473, 37.662, and 47.979 millihartree errors relative to CCSDT obtained at $\lambda=0$, 0.2, 0.4, 0.6, 0.8, and 1 with CCSD/cc-pVDZ and the associated NPE value of 21.152 millihartree to 22.183, 23.029, 23.947, 25.279, 21.842, 32.125, and 10.283 millihartree, respectively. This should be compared to the much smaller error and NPE values characterizing the corresponding CC($P$;QQ)/$c$-pVDZ calculations, which are 0.431–3.539 and 3.108 millihartree, respectively. The analogous CC($P$) calculations using $N_{\text{det(in)}}=250,000$, where the errors characterizing the $\delta(\mbox{$P$;$Q$})$-corrected CC($P$;$Q$) energies in the entire $\lambda=0$–1 region and the overall NPE relative to CCSDT are already at the level of 0.2–0.5 millihartree, produce the 17.137–18.165 millihartree errors and the NPE of 1.028 millihartree. Even with the largest $N_{\text{det(in)}}$ value considered in this study, of $10,000,000$, the 7.252–10.238 millihartree differences between the CC(PP)/$c$-pVDZ and CCSDT/cc-pVDZ energies of the lowest $^{1}A_{g}(D_{2\text{h}})$ state of cyclobutadiene in the $\lambda=0$–1 region and the NPE of 2.986 millihartree that characterizes the resulting CC($P$) potential relative to its CCSDT parent remain. All of this implies that while relaxing $T_{1}$ and $T_{2}$ amplitudes in the presence of the leading triples is important, correcting the CC($P$) energies for the remaining $T_{3}$ effects, which the CC($P$) computations using the $P$ spaces generated with the help of CIPSI do not describe, is critical to reach submillihartree accuracy levels relative to CCSDT with small fractions of triples in these spaces. We observed a similar behavior in the semi-stochastic, CIQMC- and CCMC-driven,Deustua et al. (2017); Yuwono et al. (2020); Deustua et al. (2021); Chakraborty et al. (2022) and adaptiveGururangan and Piecuch (2023) CC($P$) and CC($P$;$Q$) calculations, although based on the numerical evidence that we have generated to date, the CIPSI-driven CC($P$;$Q$) methodology investigated in this study and its recently formulated adaptive analog seem to be more effective in converging the target CC (in most of our work to date, CCSDT) energetics than their semi-stochastic counterparts. This suggests that the sequences of Hamiltonian diagonalizations utilized in the CIPSI-driven CC($P$) and CC($P$;$Q$) computations and the moment expansions defining the $\delta(\mbox{$P$;$Q$})$ corrections that are used to construct excitation spaces in the adaptive CC($P$) and CC($P$;$Q$) runs are more efficient in identifying the leading higher–than–doubly excited determinants for inclusion of the underlying $P$ spaces than the CIQMC/CCMC wave function propagations, although this topic needs to be explored further and we will return to it in the future.
@@
As shown in Table 2 and Figures 3 and 5 (c) and (d), many of the above observations apply to the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations for the lowest-energy $^{3}B_{1g}(D_{\rm 2h})$ potential, but, given the fact that the CR-CC(2,3) approach is already very accurate in this case, producing errors relative to CCSDT that in absolute value do not exceed 60 microhartree when the cc-pVDZ basis set is employed, the CC($P$;$Q$) computations using $N_{\text{det(in)}}>1$ offer no obvious advantages over CR-CC(2,3). Nonetheless, it is reassuring that, in analogy to the $A_{g}(D_{\rm 2h})$-symmetric singlet ground state, the differences between the energies of the $^{3}B_{1g}(D_{\rm 2h})$ state obtained in the CC($P$) computations using CIPSI Hamiltonian diagonalizations to create lists of the leading triply excited determinants for inclusion in the underlying $P$ spaces and their CCSDT counterparts decrease as $N_{\text{det(in)}}$ increases, independent of $\lambda$. It is also encouraging that the $\delta(\mbox{$P$;$Q$})$ corrections are as effective in bringing the CC($P$) energies to a virtually perfect agreement with the parent CCSDT data as in the case of the CIPSI-driven CC($P$;QQ)/$c$-pVDZ calculations for the lowest $^{1}A_{g}(D_{2\text{h}})$ potential using $N_{\text{det(in)}}\geq 250,000$. By inspecting the CC($P$;$Q$) column in Table 2, one may get the impression that the incorporation of the triply excited determinants identified by the CIPSI runs using increasingly large $N_{\text{det(in)}}$ values in the preceding CC($P$) steps worsens the CR-CC(2,3) results for the lowest $^{3}B_{1g}(D_{\rm 2h})$ state, which correspond to $N_{\text{det(in)}}=1$, but reading Table 2 in this way would be misleading. Indeed, in single-reference situations, such as that created by the lowest $^{3}B_{1g}(D_{\rm 2h})$ state of cyclobutadiene, where many-electron correlation effects are essentially only dynamical, the triples correction of CR-CC(2,3) [similarly to CCSD(T)] often overshoots the parent CCSDT energies (slightly). Once one starts adding triply excited determinants to the $P$ space, the CC($P$;$Q$) energies initially go up, becoming upper bounds to their CCSDT counterparts, but when the fraction of triples in the $P$ space is large enough, the differences between the CC($P$;$Q$) and CCSDT energies decrease, steadily approaching 0. We see some of this behavior in Table 2, but we have to keep in mind that the $P$ spaces used in our CIPSI-driven CC($P$;$Q$) calculations use small fractions of triples, so that the residual, $\sim$0.1 millihartree, errors relative to CCSDT remain. What is most important here is that the CC($P$;$Q$) computations for the lowest triplet potential of cyclobutadiene using $N_{\text{det(in)}}>1$ reported in Table 2 and Figures 3 and 5 (d) do not substantially alter the already excellent CR-CC(2,3) energetics. This allows us to conclude that the coupling of the lower-rank $T_{1}$ and $T_{2}$ clusters with their higher-rank $T_{3}$ counterpart is negligible in this case and the relaxation of the CCSD values of $T_{1}$ and $T_{2}$ amplitudes by including some triples in the iterative CC($P$) steps is not necessary for obtaining high-accuracy CC($P$;$Q$) results.
@@
Having demonstrated the excellent performance of the CIPSI-driven CC($P$;$Q$) approach in accurately approximating the lowest-energy $^{1}A_{g}(D_{\rm 2h})$ and $^{3}B_{1g}(D_{\rm 2h})$ potentials of cyclobutadiene along its automerization coordinate obtained with CCSDT, we now turn to the CC($P$) and CC($P$;$Q$) singlet–triplet gaps and their dependence on $\lambda$ and $N_{\text{det(in)}}$ examined, using the cc-pVDZ basis, in Table 3 and Figures 4, 5 (e) [the uncorrected CC($P$) energetics], and 5 (f) [the CC($P$;$Q$) results]. As pointed out above, to obtain accurate $\Delta E_{\text{S--T}}$ values for cyclobutadiene in the vicinity of the barrier region, one has to balance significant nondynamical correlations associated with the multiconfigurational singlet state, which manifest themselves in massive $T_{3}$ clusters that are strongly coupled to the one- and two-body components of $T$, with predominantly dynamical correlations characterizing the lowest triplet state that result in the generally smaller $T_{3}$ contributions having minimal effect on $T_{1}$ and $T_{2}$ amplitudes. The singlet–triplet gap values reported in Table 3, Figure 4, and Figure 5 (e) and (f) clearly show that neither the CCSD approach nor the CR-CC(2,3) triples correction to CCSD, which are equivalent to the CIPSI-driven CC($P$) and CC($P$;$Q$) calculations using $N_{\text{det(in)}}=1$, can do this. Both of these methods struggle with achieving a balanced description of the $^{1}A_{g}(D_{\rm 2h})$ and $^{3}B_{1g}(D_{\rm 2h})$ states of cyclobutadiene as $\lambda$ approaches 1, producing errors relative to CCSDT that in the calculations using the cc-pVDZ basis set are as large as 5.261 and 2.282 kcal/mol, respectively, at $\lambda=0.6$, where the CCSDT value of $\Delta E_{\text{S--T}}$ is 10.295-10.295 kcal/mol, 8.589 and 4.434 kcal/mol, respectively, at $\lambda=0.8$, where $\Delta E_{\text{S--T}}$ obtained with CCSDT is 5.804-5.804 kcal/mol, and 15.120 and 9.222 kcal/mol at $\lambda=1$, where the CCSDT result for $\Delta E_{\text{S--T}}$ is 4.783-4.783 kcal/mol. These large error values in the singlet–triplet gaps resulting from the CCSD and CR-CC(2,3) computations in the barrier region are a consequence of the dramatic increase in the magnitude of $T_{3}$ effects characterizing the $^{1}A_{g}(D_{\rm 2h})$ state and a rapidly deteriorating description of this state by both CCSD and CR-CC(2,3) as $\lambda\rightarrow 1$, seen in Table 1, Figure 2, and Figure 5 (a) and (b), as opposed to the nearly constant $T_{3}$ contributions and a virtually perfect agreement between the CR-CC(2,3) and CCSDT $^{3}B_{1g}(D_{\rm 2h})$ potentials in the entire $\lambda=0$–1 region shown in Table 2, Figure 3, and Figure 5 (d) [while quantitatively inaccurate, the shape of the $^{3}B_{1g}(D_{\rm 2h})$ potential obtained with CCSD, shown in Figure 5 (c), is qualitatively correct too]. In analogy to the previously discussed calculations for the lowest-energy $^{1}A_{g}(D_{\rm 2h})$ potential, in order to bring the above errors down, the input parameter $N_{\text{det(in)}}$, which controls the CIPSI diagonalization sequences preceding the CC($P$) and CC($P$;$Q$) computations, must be increased, so that the $P$ spaces used in these computations are augmented with the leading triply excited determinants, the CCSD values of $T_{1}$ and $T_{2}$ amplitudes, used in CR-CC(2,3), are properly relaxed, and the quality of the $\delta(\mbox{$P$;$Q$})$ corrections that capture the remaining $T_{3}$ effects improves.